Circular No. 8307

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004av IN ESO 571-G15 Further to IAUC 8305, M. Moore and W. Li report the LOSS discov-

ery, on KAIT images taken on Mar. 17.35 and 18.35 UT, of an apparent supernova (mag ∼ 15.8) located at α = 11^{h}40^{m}57^{s}.84, δ = −22^{o}28′53′′.6 (equinox 2000.0), which is 12′′.6 west and 15′′.9 south of the nucleus of ESO 571-G15. A KAIT image taken on Mar. 4.32 also showed the new object at mag ∼ 16.8, while an image taken on Jan. 22.36 showed nothing at this position (limiting mag ∼ 18.0).

Further to IAUC 8303, T. Matheson et al. report that a spectrum of SN 2004av, obtained by P. Berlind on Mar. 19.31 UT, shows it to be a type-Ia supernova a few days after maximum. Adopting the NED recession velocity of 7057 km/s for the host galaxy, the supernova expansion velocity is 9500 km/s for Si II (rest 635.5 nm). The spectral-feature age of the supernova is 4 ± 2 days after maximum light.

V5114 SAGITTARII = NOVA SAGITTARII 2004 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V5114 Sgr has been given to this nova (cf. IAUC 8306). H. Sato and T. Nakamura, National Astronomical Observatory, Tokyo, report that an independent discovery of this nova (at mag 8.9) was made by Yuji Nakamura (Mie prefecture, Japan) on Tri-X film exposed on Mar. 15.816 UT with a 200-mm f/4 camera lens, the position given as α = 18^{h}19^{m}30^{s}.4, δ = −28^{o}36′41′′ (equinox 2000.0; uncertainty ± 30′′). H. Yamaoka, University of Kyushu, reports that V5114 Sgr was not visible on ASAS-3 V images taken on Mar. 11.37 (limiting mag ∼ 13.5).

M. Della Valle, Istituto Nazionale di Astrofisica, Arcetri, Florence; and A. Ederoclite, L. Schmidtobreick, L. Germany, T. Dall, and I. Saviane, European Southern Observatory, write: “Preliminary analysis of a spectrum (range 380–900 nm, resolution 48000), obtained on Mar. 18.3 UT at La Silla with the 2.2-m telescope (+ Fiber-fed Extended Range Optical Spectrograph), shows that V5114 Sgr is indeed a nova caught near maximum light. The spectrum is dominated by H lines, Ca II (mult. 2), and O I (mult. 34, 1) and Na I emission lines. Fe II emission lines (mult. 74, 55, 49, 42) are also detected, although they are still weak. All Balmer lines are flanked by double P-Cyg profiles whose minimum absorptions give maximum expansion velocities of ∼ 1500 and 800 km/s, respectively. The equivalent widths of the Na D interstellar lines (0.11 and 0.10 nm) may suggest that this nova suffers from moderate absorption.”

2004 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green